A mixture contains 1.0 x 10-3 M Cu+2 and Fe+2 cations, and is saturated with .10 M H2S. Determine the pH where CuS precipitates, but FeS does not precipitate. Ksp for CuS = 4.0 x 10-36 and Ksp for FeS = 4.0 x 10-17
Answer #1

CuS    ⇌   Cu 2+ +S 2-

Ksp for CuS= 4*10-36

FeS      ⇌          Fe 2+ +S 2-

Ksp for FeS=4*10-17  

H2S    ⇌            2H+ +S 2-

The sulfide concentrations needed to cause CuS and FeS to precipitate are

Ksp=[Cu2+] [S2−]

4*10-36 = (0.10M) [S2−]

[S2−]=4*10-36/1*10 -1=4*10 -35 M       (1)

and

Ksp=[Fe 2+][S2-]

4*10-17   =(0.10M) [S2−]

[S2−]=4*10 -16 M      (2)

sulfide concentrations between 4*10 -35 M and 4*10 -16 M      will precipitate CuS and not FeS.

H2S       ⇌     2H+ +S 2-

pKa1=6.9=-logKa1

Ka1=10-pKa = 1.1*10-7   (3)

         

HS    ⇌    H++S2-

pKa2=12.9

Ka2=1.3×1013       (4)

The pKa1for H2S is 6.97, and pKa2 corresponding to the formation of [S2−] is 12.90

Ka1=1.1*10-7   = [H+][HS] / [H2S]

= x2 / 0.10 M

x =1.1×10−4 M

= [H+]=[HS]

Substituting this value for [H+] and [HS] into (4)

Ka2=1.3×1013= [H+][S2−] / [HS−]=(1.1×104 ) x /1.1×104

x=[S2]

equilibrium constant K for the overall reaction, which is the product of Ka1 and Ka2, and the concentration of H2S in a saturated solution to calculate the H+ concentration needed to produce [S2−] of

K=Ka1 Ka2

= (1.1×10−7)(1.3×10−13)=1.4×10−20   = [H+][HS] / [H2S]* [H+][S2−] / [HS−]

=[H+][H+][S2-] / [H2S]

= [H+]2 [S2-] / [H2S]

[H+]2=K[H2S][S2-]=(1.4×10−20)(0.10 M) / 4*10 -35 M

[H+] = 0.35*10 14

pH = -log10[H+]

       =-log10 0.35*10 14

= -log3.5*1013

=12.45